# Linear Projections of Teacher Embeddings for Few-Class Distillation

### v1

@@ 1 Introduction
    Motivated by recent insights into the Neural Collapse phenomenon, we demonstrate that the invention of pseudo-classes through unsupervised clustering of teacher embeddings can enhance distillation performance in binary and few-class classification tasks. However, the efficacy of this approach is contingent upon the specific clustering algorithm employed.
@@ 1.1 Organization of the paper
In Section 2 we discuss related work, including the baselines we considered in this work and the Neural Collapse phenomenon. In Section 3 we present our method. In Section 4 we present our experiments. Specifically, in Section 4.2 we present experiments demonstrating that inventing pseudo-classes via clustering of the teacher’s model’s embeddings can improve distillation effectiveness, and in Section 4.3 we compare LELP with other distillation baselines. In Appendix A we discuss the broader impact of our work. In Appendix B we present our experimental results corresponding to binary classification tasks with subclass structure in detail. In Appendix C we provide ablations on the design choices behind LELP. In Appendix D we give a detailed description of the NLP datasets we used. In Appendix E we present results regarding the applicability of LELP to multiclass classification tasks with a moderate number of classes. In Appendix F we study additional properties of LELP, including its relative performance in the semi-supervised setting [7, 30, 15, 4, 17], its data efficiency and its training speed. In Appendix G we provide pseudo-code for our method. Finally, in Appendix H we provide implementation details.
@@ 4 Experimental Evaluation
In this section we present our experimental results. In Section 4.1 we describe our experimental setup. Section 4.2 presents experiments showcasing that generating pseudo-classes via unsupervised clustering of the teacher model’s embeddings can improve distillation effectiveness. In Section 4.3 we compare LELP with other distillation baselines.
@@ 4.1 The Setup
Figure 3: The effectiveness of different clustering techniques for creating pseudo-classes during knowledge distillation from a ResNet-92 teacher to (a) a ResNet-56 and (b) a MobileNet student on the binarized CIFAR-100 dataset is presented. “Oracle Clustering”, where subclass structure is known a priori, serves as an upper bound and notably surpasses all other methods, even exceeding the teacher’s performance in the ResNet-56 case. Among practical methods (i.e., those discovering subclass structure), LELP exhibits superior performance. Agglomerative and K-means clustering do not consistently outperform vanilla knowledge distillation, demonstrating the dependence of pseudo-class effectiveness on the chosen clustering algorithm.
@@
4.2 Investigating methods for inventing pseudo-classes
@@
In this section we investigate the extent to which inventing pseudo-classes via unsupervised clustering of the teacher model’s embeddings can be beneficial for distillation. Our investigation focuses on tasks with inherent subclass structure, driven by the hypothesis that class embeddings contain rich and informative structural details. Specifically, we utilize binarized versions of CIFAR-10 and CIFAR-100, assigning binary labels based on the original class labels ($y_{binary}=y_{original}\%2$). These datasets present a unique challenge for knowledge distillation, as vanilla distillation is known to underperform due to limited class label information.
@@
Crucially, the availability of original labels ($y_{original}$) in these subclass datasets allows us to explore the "Oracle Clustering" approach. This entails training the student model on the full 10/100-way classification task for CIFAR-10/100, respectively, and subsequently treating the problem as binary classification during testing. Interestingly, the student models trained utilizing the Oracle Clustering approach exhibit the highest performance among all evaluated clustering methods. This approach not only outperforms other clustering strategies but also surpasses the performance of the original teacher network in certain cases. A t-SNE visualization of the feature embeddings learned by the student models, presented in Figure 7, offers a deeper understanding of the underlying factors contributing to this observed performance advantage. While Oracle Clustering represents an idealized scenario where subclass structure is known a priori and thus impractical for real-world datasets, its consideration underscores the potential power of inventing pseudo-classes as a methodological approach.
@@
We consider several natural ways of inventing pseudo-classes by clustering the teacher’s embedding space and we compare them with LELP. In particular, we consider three approaches: (i) Agglomerative clustering (complete linkage); (ii) K-means clustering; and (iii) K-means clustering after we first have projected the teacher’s embedding space in a two-dimensional space using t- SNE [13] as proposed in [34]. For these methods, we tested both hard-label (one-hot vector) and soft-label (weighted by teacher probabilities) representations of the subclasses during student training. We observed that soft-labels, even with temperature tuning, did not significantly impact performance. Therefore, we present results using only hard-label (one-hot vector) subclass representations, which can be found in Table 1 (see also Figure 3). Details on the number of clusters chosen can be found in Appendix H.3.
@@
Analysis of Table 1 reveals several key findings: (i) LELP consistently outperforms all other clustering methods as well as Vanilla KD across all experimental scenarios; (ii) t-SNE & K-means generally demonstrate superior performance to Vanilla KD; (iii) Agglomerative clustering and K-means clustering exhibit more varied results in comparison to Vanilla KD. Thus, our findings suggest that the generation of pseudo-classes via clustering methodologies holds significant potential, however, the efficacy of this approach is contingent upon the specific clustering algorithm employed.
@@ 6 Conclusion
In this study, we have presented evidence that the creation of pseudo-classes via unsupervised clustering of teacher embeddings can improve distillation performance in binary and few-class classification tasks, without necessitating the retraining of the teacher model. Through empirical evaluation, we observed that linear projections consistently yield high performance, prompting us to introduce LELP. The superior performance of the "Oracle Clustering" method, where subclass structure is known a priori, suggests that the generation of pseudo-classes through clustering techniques has substantial promise. Consequently, future research can explore more sophisticated methods for extracting teacher embedding information, drawing insights from the expanding body of work on Neural Collapse, or investigate strategies for distilling intermediate layer embeddings using a similar approach.

### v2

@@ 1 Introduction
    Motivated by recent insights into the Neural Collapse phenomenon, we demonstrate that the invention of pseudo-subclasses through unsupervised clustering of teacher embeddings can enhance distillation performance in binary and few-class classification tasks. However, the efficacy of this approach is contingent upon the specific clustering algorithm employed.
@@ 1.1 Organization of the paper
In Section 2 we discuss related work, including the baselines we considered in this work and the Neural Collapse phenomenon. In Section 3 we present our method. In Section 4 we present our experiments. Specifically, in Section 4.2 we present experiments demonstrating that inventing pseudo-subclasses via clustering of the teacher’s model’s embeddings can improve distillation effectiveness, and in Section 4.3 we compare LELP with other distillation baselines. In Appendix A we discuss the broader impact of our work. In Appendix B we present our experimental results corresponding to binary classification tasks with subclass structure in detail. In Appendix C we provide ablations on the design choices behind LELP. In Appendix D we give a detailed description of the NLP datasets we used. In Appendix E we present results regarding the applicability of LELP to multiclass classification tasks with a moderate number of classes. In Appendix F we study additional properties of LELP, including its relative performance in the semi-supervised setting [7, 30, 15, 4, 17], its data efficiency and its training speed. In Appendix G we provide pseudo-code for our method. Finally, in Appendix H we provide implementation details.
@@ 4 Experimental Evaluation
In this section we present our experimental results. In Section 4.1 we describe our experimental setup. Section 4.2 presents experiments showcasing that generating pseudo-subclasses via unsupervised clustering of the teacher model’s embeddings can improve distillation effectiveness. In Section 4.3 we compare LELP with other distillation baselines.
@@ 4.1 The Setup
Figure 3: The effectiveness of different clustering techniques for creating pseudo-subclasses during knowledge distillation from a ResNet-92 teacher to (a) a ResNet-56 and (b) a MobileNet student on the binarized CIFAR-100 dataset is presented. “Oracle Clustering”, where subclass structure is known a priori, serves as an upper bound and notably surpasses all other methods, even exceeding the teacher’s performance in the ResNet-56 case. Among practical methods (i.e., those discovering subclass structure), LELP exhibits superior performance. Agglomerative and K-means clustering do not consistently outperform vanilla knowledge distillation, demonstrating the dependence of pseudo-subclass effectiveness on the chosen clustering algorithm.
@@
4.2 Investigating methods for inventing pseudo-subclasses
@@
In this section we investigate the extent to which inventing pseudo-subclasses via unsupervised clustering of the teacher model’s embeddings can be beneficial for distillation. Our investigation focuses on tasks with inherent subclass structure, driven by the hypothesis that class embeddings contain rich and informative structural details. Specifically, we utilize binarized versions of CIFAR-10 and CIFAR-100, assigning binary labels based on the original class labels ($y_{binary}=y_{original}\%2$). These datasets present a unique challenge for knowledge distillation, as vanilla distillation is known to underperform due to limited class label information.
@@
Crucially, the availability of original labels ($y_{original}$) in these subclass datasets allows us to explore the "Oracle Clustering" approach. This entails training the student model on the full 10/100-way classification task for CIFAR-10/100, respectively, and subsequently treating the problem as binary classification during testing. Interestingly, the student models trained utilizing the Oracle Clustering approach exhibit the highest performance among all evaluated clustering methods. This approach not only outperforms other clustering strategies but also surpasses the performance of the original teacher network in certain cases. A t-SNE visualization of the feature embeddings learned by the student models, presented in Figure 7, offers a deeper understanding of the underlying factors contributing to this observed performance advantage. While Oracle Clustering represents an idealized scenario where subclass structure is known a priori and thus impractical for real-world datasets, its consideration underscores the potential power of inventing pseudo-subclasses as a methodological approach.
@@
We consider several natural ways of inventing pseudo-subclasses by clustering the teacher’s embedding space and we compare them with LELP. In particular, we consider three approaches: (i) Agglomerative clustering (complete linkage); (ii) K-means clustering; and (iii) K-means clustering after we first have projected the teacher’s embedding space in a two-dimensional space using t- SNE [13] as proposed in [34]. For these methods, we tested both hard-label (one-hot vector) and soft-label (weighted by teacher probabilities) representations of the subclasses during student training. We observed that soft-labels, even with temperature tuning, did not significantly impact performance. Therefore, we present results using only hard-label (one-hot vector) subclass representations, which can be found in Table 1 (see also Figure 3).
@@
Analysis of Table 1 reveals several key findings: (i) LELP consistently outperforms all other clustering methods as well as Vanilla KD across all experimental scenarios; (ii) t-SNE & K-means generally demonstrate superior performance to Vanilla KD; (iii) Agglomerative clustering and K-means clustering exhibit more varied results in comparison to Vanilla KD. Thus, our findings suggest that the generation of pseudo-subclasses via clustering methodologies holds significant potential, however, the efficacy of this approach is contingent upon the specific clustering algorithm employed.
@@ 6 Conclusion
In this study, we have presented evidence that the creation of pseudo-subclasses via unsupervised clustering of teacher embeddings can improve distillation performance in few-class classification tasks, without necessitating the retraining of the teacher model. Through empirical evaluation, we observed that linear projections consistently yield high performance, prompting us to introduce LELP. The superior performance of the "Oracle Clustering" method, where subclass structure is known a priori, suggests that the generation of pseudo-subclasses through clustering techniques has substantial promise. Consequently, future research can explore more sophisticated methods for extracting teacher embedding information, drawing insights from the expanding body of work on Neural Collapse, or investigate strategies for distilling intermediate layer embeddings using a similar approach.